Grade 12Physical Chemistry

A mixture has 18 gram water and 414 gram ethanol the mole fraction of water in mixture is (assume ideal behaviour of the mixture) ?

Ashish Pardeshi
8 Years agoGrade 12

2 Answers

Malay Damani
8 Years ago
number of moles of water = 18/18=1 mol
number of moles of ethanol=414/46=9
mole fraction of water=1/(1+9)=1/10=0.1
Shivam kumar
7 Years ago
Mole of H2O = 1 
Mole.of ethanol =9
Total mole present in mixture =10 mole 
Mole fraction of water =1/10=0.1 mole